Institut for Matematik og Datalogi Odense Universitet March 23, 1995

JFB

DM15 – F95 – Week 23

Lecture, March 20

We finished chapter 11 in Sudkamp, and section 12.1 of chapter 12.

Lecture, March 27

We will finish chapter 12 in Sudkamp, and begin on chapter 13.

Lecture, April 3

We will finish chapter 13 in Sudkamp, and begin on the first chapter of Introduction to Programming Language Theory, by John C. Mitchell. This is available in the bookstore for 39 kr.

Problems to be discussed in week 14

• Do problem 2b on page 296 of Sudkamp.

• Do problem 3b on page 297 of Sudkamp.

• Do problem 6c on page 297 of Sudkamp.

• Do problem 7 on page 297 of Sudkamp.

• Do problem 10 on page 298 of Sudkamp.

• Do problem 11 on page 298 of Sudkamp.

• Do problem 12 on page 298 of Sudkamp.

• Do problem 14 on page 298 of Sudkamp.